OpenStudy (anonymous):

Find the 2 points where the curve x^2+xy+y^2=7 crosses the x-axis, and show that the tangents to the curve at these points are parallel. what is the common slope of these tangents? No clue. Help is much appreciated :)

OpenStudy (anonymous):

wherever the curve crosses x-axis, you cn use the condition y=0

OpenStudy (anonymous):

@him1618 then what?

OpenStudy (anonymous):

cn u tell me the eqn now?

OpenStudy (anonymous):

x^2=7?

OpenStudy (anonymous):

put y=0 in ur eqn u get x^2 = 7

OpenStudy (anonymous):

solve it for

OpenStudy (anonymous):

x=7^1/2?

OpenStudy (anonymous):

+ root 7 and - root 7

OpenStudy (anonymous):

that's it?

OpenStudy (anonymous):

so the points are (+root7, 0) and (-root7, 0)

OpenStudy (anonymous):

yes, that's it?

OpenStudy (anonymous):

now u need to find the slopes at these two points

OpenStudy (anonymous):

can u find y' from ur eqn?

OpenStudy (anonymous):

thank you :)